Required information The Foundational 15 (LO4-1, LO4-2, LO4-3, L04-4, L04-5] [The following information applies to the questi8. What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.) Cost per equivalent unit for9. What is the cost per equivalent unit for conversion? (Round your answer to 2 decimal places.) Cost per equivalent unit10. What is the cost of ending work in process inventory for materials? (Round your intermediate calculations to 2 places.) C11. What is the cost of ending work in process inventory for conversion? (Round your intermediate calculations to 2 places.)12. What is the cost of materials transferred to finished goods? (Round your intermediate calculations to 2 places.) Cost of
Answer #1

WN-1

Statement of Equivalent Production

Materials CC**     

Particulars Quantity Quantity % Quantity %
1.Output (FG)* 34500 34500 100 34500 100
2.Closing WIP 9000 9000 100 9000×50%=4500 50
3.Equivalent Production(1+2) 43500 43500 39000

Here OH means Overheads,FG means Finished Goods.

*Output=Opening WIP+Material Added-Closing WIP

=5500u+38000u-9000u=34500u

CC** means conversion, which is labor and overhead.

WN-2

Statement of Cost per Equivalent unit

Particulars Material Conversion Cost(Labor+Overhead)
1.Cost 128325 186500(84500+102000)
2.Opening WIP 21750 18250
3.Total 150075 204750
4.Equivalent Production units(from WN-1) 43500 39000
5.Cost per Equivalent unit($)[3÷4] 3.45 5.25

Total Cost per unit = $(3.45+5.25)=$8.7

(8)Cost per Equivalent unit for materials(from WN-2) =$3.45

(9)Cost per Equivalent unit for Conversion (from WN-2)=$5.25

The below calculations are done on the basis of WN-1,WN-2

(10)Cost of ending work in progress inventory =9000u×$3.45=$31050

(11)Cost of ending work in progress for conversion = 4500u×$5.25=$23625

(12) Cost of materials transferred to Finished goods = 34500u×$8.7=$300150